# HBS - Hardware Build System: A Tcl-based, minimal common abstraction approach for build system for hardware designs

### v3

@@ Abstract.
Build systems become an indispensable part of the software implementation and deployment process. New programming languages are released with the build system integrated into the language tools, for example, Go, Rust, or Zig. However, in the hardware description domain, no official build systems have been released with the predominant Hardware Description Languages (HDL) such as VHDL or SystemVerilog. Moreover, hardware design projects are often multilanguage.
The paper proposes a new build system for the hardware description domain. The system is called the Hardware Build System (HBS). The main goals of the system include simplicity, readability, a minimal number of dependencies, and ease of integration with the existing Electronic Design Automation (EDA) tools. The system proposes a novel, minimal common abstraction approach, whose particular implications are described in the article. All the core functionalities are implemented in Tcl. Only the EDA tool’s independent features, such as dependency graph generation, are implemented in a Python wrapper.
@@ 1. Introduction
Implementation of any complex software or hardware description design involves not only code writing. Equally important is defining metadata used for configuration, dependency management, code generation, testing, and deployment purposes. The metadata is read by software tools, which, in turn, are responsible for generating final binary or bitstream files, or for carrying out extra tasks such as unit tests. Such programs are often referred to as build systems. Some build systems are split into multiple smaller programs, and some have a monolithic structure. An introduction to the build system concept and principles can be found in (Smith, 2011; MOKHOV et al., 2020).
@@
In the software domain, there are multiple well-established build systems, such as GNU Make (Free Software Foundation, 2025), CMake (CMake, 2025), Ninja (Martin, 2025), or Bazel (Google, 2025). Some of them were implemented in the early days of programming. For example, GNU Make was first released in 1988 as part of the GNU Project. It was implemented to provide a free software alternative to Unix Make, which was first implemented in 1976 by Stuart Feldman at Bell Labs. Modern programming languages, such as Go, Rust, or Zig, come with build systems included in the language toolchains. The Zig build system can even be used as a build system for C/C++ projects.
@@
In the hardware domain, the situation looks different. The earliest Electronic Design Automation (EDA) tools were implemented as proprietary programs. EDA vendors did not (and still do not) cooperate to create a standard (or at least similar) user interface for tool interaction. The main common point between various EDA tools is using the Tcl language as the interpreter for interaction with the tools. However, the tools provide custom commands and parameters even for primary, common tasks, such as adding hardware description files to a project design. For example, to add a VHDL file in AMD Vivado, one can use specific read_vhdl <path_to_vhdl_file> or generic add_files <path_to_vhdl_file> commands. To add a VHDL file in Intel Quartus, one has to use set_global_assignment -name VHDL_FILE <path_to_vhdl_file> command. EDA tools have abundant differences, making implementing a build system for hardware projects harder than for software. Despite this, there were (and still are) attempts to implement build systems that are more suitable for hardware projects.
The paper proposes a new build system for a hardware description domain. The system is Tcl-based, which makes it easy to integrate with all existing EDA tools already utilizing Tcl language as a user-tool interface. A novel, minimal common abstraction approach differentiates the implemented system from the existing build systems targeting the hardware domain.
The structure of the article is as follows. Section 2 describes related work. Section 3 discusses the requirements and their reasons imposed on the implemented build system. Section 4 introduces the internal architecture of the implemented build system. Section 5 describes commands provided by the implemented build system. Section 6 evaluates the proposed solution in terms of other build systems. Section 7 summarizes the work and highlights novel aspects of the implemented build system.
2. Related work
@@
There already exist at least more than ten build systems targeting the hardware domain. Some of them are vendor tool specific, and some try to be generic, entirely abstracting away the tool used for synthesis or implementation. By looking at the architecture (analyzing how a tool internally works), existing hardware build systems can be classified into two approaches:
@@
    direct Tcl approach,
@@
    indirect abstract approach.
2.1. Direct Tcl approach
In the Tcl approach, the user directly utilizes the Tcl interface. The idea seems obvious. Since EDA tools already utilize Tcl as a user interface, a build system should also provide users with a Tcl interface. Usually, the user does not directly call EDA tool specific commads. Tcl approch build systems implement some wrapper commands that the user should use. Representatives of the direct Tcl approach are, for example, fpga-vbs (Braun, 2025), vextproj (Zabołotny, 2025, 2016), and OSVVM-Scripts (Lewis, 2025).
Being able to easily execute arbitrary, custom EDA tool commands is the most significant advantage of the direct Tcl approach. As a Tcl interpreter embedded into an EDA tool executes the build system code, a user can call custom commands in almost any place.
The direct Tcl approach is not free of drawbacks. Some of them are inherent, and some result from the fact that direct Tcl build systems development was stopped at some stage by their maintainers.
The first inherent drawback is that Tcl is considered an unfriendly language compared to, for example, Python. The argument is often raised by junior hardware engineers and has a very subjective nature. However, Tcl was chosen as the EDA scripting language because it was very popular when EDA tools were developed (Chen et al., 2001). The second inherent drawback is that Tcl is not a popular language nowadays. Although there are no physical obstacles to change this trend, it is doubtful that suddenly more developers start using Tcl instead, for example, Python as an interpreted language. This is why this drawback is also treated as inherent.
The second group of direct Tcl build systems drawbacks results from stopping the development at a certain stage. There is no direct Tcl build system that can be used to examine the limits and bottlenecks of this idea. For example, most of the direct Tcl build systems support only one EDA tool. OSVVM-Scripts is an exception. However, it can be used only for simulation. It does not support any synthesis or implementation tools. More example drawbacks common to the direct Tcl build systems include:
@@ 2.1. Direct Tcl approach
    Lack or limited capability of automatic testbench targets detection.
@@
2.2. Indirect abstract approach
The indirect abstract approach is much more popular than the direct Tcl approach. In this approach, the user declares various properties and configurations instead of directly writing Tcl code or calling EDA tools Tcl commands.
The most popular hardware build system utlizing the indirect abstract approach is probably FuseSoc (Kindgren, 2019, 2025b). Other popular build systems include SiliconCompiler (Olofsson et al., 2022; sil, 2025), HAMMER (Liew et al., 2022; ham, 2025), Hog (Biesuz et al., 2021; Gonnella et al., 2025), hdlmake (hdl, 2025), orbit (Ruskin, 2025), bender (ben, 2025), bazel_rules_hdl (baz, 2025), flgen (Ishitani, 2025), and Blockwork (blo, 2025).
The language and syntax used for declarations vary between different build systems. Some use simple data serialization formats such as YAML (for example, FuseSoc or bender) or TOML (for example, orbit). Some come with a domain-specific language (for example, flgen), or use Python (for example, HAMMER) and Makefiles (for example, Hdlmake). Finally, some use a mix of more than one (for example, Hog). Especially using Python as a declarative language is powerful, as users can use all functionalities that come with the language.
The most significant advantage of the indirect abstract approach is the low entry threshold. The syntax of languages such as YAML, TOML, or Python is simple, which leads to high legibility. Even junior engineers without extra training can understand it. Listing 1 presents a very basic core definition in the FuseSoc build system. The first line specifies the version of the description format. The “name” key specifies the vendor, library, name, and version (VLNV) identifier specified in the IP-XACT User Guide (Accellera IP-XACT Working Group, 2018). The remaining part is relatively self-explanatory.
@@ 2.2. Indirect abstract approach
The indirect abstract approach is based on the concept that the interface to EDA tools is hidden under some common abstraction layer. Whether this is an advantage or a disadvantage is debatable. Providing a common abstract layer means that users do not have to learn the interface of multiple EDA tools. Moreover, in theory, changing the EDA tool for compilation process should be limited to changing solely the EDA tool declaration. However, this is somewhat a utopian view. As practice shows the common abstraction layer leads to the following common difficulties:
@@
    Missing feature implementation for some EDA tools. Let us suppose the abstraction layer declares feature F, and the build system claims support for tools T1 and T2. Although T1 might provide the implementation for F, T2 might be missing implementation for this particular feature. In such a case, the user must add support for this feature to the build system. This task might not be trivial. The user must get to know the internals of the build system, prepare a pull request, wait for a review, and potentially reiterate the implementation. Yet another approach is to modify automatically generated build scripts before a compilation. However, this is against the whole idea of build systems. A direct Tcl approach is free of these difficulties because the user can directly insert custom Tcl commands. There are no intermediate abstraction layers.
@@
All indirect abstract hardware build systems allow “injecting” arbitrary custom Tcl commands into the automatically generated scripts. However, these commands can not be inserted into arbitrary places. Moreover, inserting a single-line Tcl command usually requires more than one line of code or even more than one file. This can be treated as a proof that the indirect abstract approach is not a golden bullet. The idea of a common abstraction layer has significant drawbacks that no one yet discovered how to fix.
Making a truly abstract hardware build system is highlighy unliekly in the nearest feature. The main problem with the common abstraction layer is the fact that EDA tools were implemented before such a layer was proposed. Declaring a common interface for already existing tools is simply much harder task than implementing tools conforming to some formal interface.
The HBS described in the paper proposes a novel approach. It tries to incorporate the best of the direct Tcl approach and indirect abstract approach. It implements a common abstraction layer. However, the layer is:
@@
    limited to the primary common features,
@@
    implemented in Tcl.
@@ 3.1. Simplicity and ease of use
The primary goal of the hardware build system proposed in the paper is simplicity. Contemporary hardware build systems utilizing the indirect abstract approach offer simplicity only at the initial stage. However, as the project complexity grows, so does the complexity of the build system utilization.
@@ 3.2. Minimal user interface
Xia et al. in (Xia et al., 2013) report that even up to 21.35% of the bugs in software build systems belong to the external interface category. This means that one-fifth of build system bugs results from incorrect usage of the user interface or different behavior of the build system with different tools or platforms.
Fixing build system external interface bugs requires applying patches in the build system itself. Such a task is relatively time-consuming and might not be simple to implement. Build system user must report the problem to the build system developers or fix the bug himself. However, this implies that the build system user is distracted from his own work. Moreover, the fix is not applied immediately. This means that the user must also find a temporary workaround. This, in turn, leads to a technical debt (Besker et al., 2019, 2018).
A minimal user interface has one more advantage. It allows developers to customize build systems to the project needs. Each project is in some way unique and might require particular approach for build problems solving. Moreover, projects usually evolve, for example, due to the customer demand. Dynamic tailoring to the project requirements is a desired feature of modern build systems (Maudoux and Mens, 2018). Similar observations were made by Neitsh et al. in (Neitsch et al., 2012). Authors report that concerns from the application and implementation domains may “leak” into the build model. The build system should be adaptable enough to provide ways to respond to these concerns. Too rigid user interfaces limit such capabilities.
@@ 3.3. Low number of dependencies
Xiao et al. in (Xiao et al., 2022) report that dependency management is one of the top three factors of technical debt introduction in build systems. Such a finding is not surprising. Dependencies often break in unexpected moments. In such a case, long-term quality is sacrificed for the short-term goals. Getting a project to compile successfully is more important than fixing broken build system dependencies.
One of the goals of the HBS proposed in the paper is to minimize the probability of user distraction caused by broken dependencies. This implies that the number of dependencies should be kept as low as possible.

### v4

# HBS - Hardware Build System: Characterizing and comparing direct-Tcl and indirect-abstract approaches for hardware build systems

@@ Abstract.
Build systems become an indispensable part of the software implementation and deployment process. New programming languages are released with the build system integrated into the language tools, for example, Go, Rust, or Zig. However, in the hardware description domain, no official build systems have been released with the predominant Hardware Description Languages (HDL) such as VHDL or SystemVerilog. Moreover, hardware design projects are often multilingual.
The paper characterizes and compares two common approaches for hardware build system implementations. The first one, the direct-Tcl approach, in which the build system code is executed directly by the EDA tool during the design build flow. The second one, the indirect-abstract approach, in which the build system produces a Tcl script, which is later run by a proper EDA tool. As none of the existing direct-Tcl build systems was close to the indirect-abstract build systems in terms of supported functionalities, the paper also presents a new direct-Tcl hardware build system called HBS. The implemented build system was used as a representative of direct-Tcl build systems in comparison with indirect-abstract build systems.
@@ 1. Introduction
Implementation of any complex software or hardware description design involves not only code writing. Equally important is defining metadata used for configuration, dependency management, code generation, testing, and deployment purposes. The metadata is read by software tools, which, in turn, generate final binary or bitstream files or perform additional tasks, such as unit tests. Such programs are often referred to as build systems. Some build systems are split into multiple smaller programs, and some have a monolithic structure. An introduction to the build system concept and principles can be found in (Smith, 2011; MOKHOV et al., 2020).
@@
In the software domain, there are multiple well-established build systems, such as GNU Make (Free Software Foundation, 2025), CMake (CMake, 2025), Ninja (Martin, 2025), or Bazel (Mithro and others, 2021; Google, 2025). Some of them were implemented in the early days of programming. For example, GNU Make was first released in 1988 as part of the GNU Project. It was implemented to provide a free software alternative to Unix Make, which was first implemented in 1976 by Stuart Feldman at Bell Labs (Feldman, 1979). Modern programming languages, such as Go, Rust, or Zig, include build systems in their toolchains. The Zig build system can even be used as a build system for C/C++ projects.
@@
In the hardware domain, the situation looks different. The earliest Electronic Design Automation (EDA) tools were implemented as proprietary programs. EDA vendors did not (and still do not) cooperate to create a standard (or at least similar) user interface for tool interaction. The main common point among various EDA tools is the use of the Tcl language as the interpreter for interacting with the tools. However, the tools provide custom commands and parameters even for primary, common tasks, such as adding hardware description files to a project design. For example, to add a VHDL file in AMD Vivado, one can use specific read_vhdl <path_to_vhdl_file> or generic add_files <path_to_vhdl_file> commands. To add a VHDL file in Intel/Altera Quartus, one has to use set_global_assignment -name VHDL_FILE <path_to_vhdl_file> command. EDA tools have abundant differences, making implementing a build system for hardware projects harder than for software. Despite this, there have been (and still are) attempts to implement build systems better suited to hardware projects.
The main goal of the paper is to characterize and compare two existing approaches for hardware build system implementation, the direct-Tcl and indirect-abstract. The paper also partially describes the new hardware build system (HBS), which follows the direct-Tcl approach. The system was designed and implemented to serve as a comparison representative.
The article’s structure is as follows. Section 2 describes two commonly used approaches for implementing hardware build systems. Section 3 explains why HBS was implemented. Section 4 describes requirements imposed on the HBS implementation and user interface. Section 5 briefly introduces the internal architecture of HBS. Section 6 compares various aspects of the direct-Tcl and indirect-abstract approaches. Section 7 summarizes the work.
2. Hardware build system approaches
@@
There exist at least more than ten build systems targeting the hardware domain. Some of them are vendor-tool-specific, while others aim to be generic, entirely abstracting away the tool used for synthesis or implementation. By looking at the architecture (analyzing how a tool internally works), existing hardware build systems can be classified into two approaches:
@@
    Direct-Tcl approach - the build system code is executed directly by the EDA tool during the design build flow. The build system code and the build code reside in the same domain. The build system code has access to the build code and vice versa.
@@
    Indirect abstract approach - the build system code produces a Tcl script, which is only later run by a proper EDA tool. The build system code and the build code reside in different domains. The build system code generates the build code, but it does not have access to it during the design build flow. The build code is not aware of the build system code and has no access to it.
2.1. Direct-Tcl approach
In the direct-Tcl approach, the user directly writes Tcl code and uses the EDA tools’ Tcl interface. The idea seems obvious. Since EDA tools already utilize Tcl as a user interface, a build system should also provide users with a Tcl interface.
Usually, the direct-Tcl approach build systems implement some wrapper commands that the user should use for primary tasks, such as, file addition, library setting, or generic/parameter setting. However, in the case of more specific tasks, the user is supposed to call the EDA tool’s specific commands explicitly. It is also up to the user to maintain build compatibility across various EDA tools when build requirements are more complex, for example, when constraints depend on synthesis results. It is often achieved with tool guard clauses. For example, an if statement block calling a specific set of commands only if the currently executed EDA tool is specific. Representatives of the direct-Tcl approach are, for example, fpga-vbs (Braun, 2025), vextproj (Zabołotny, 2025, 2016), and OSVVM-Scripts (Lewis, 2025).
Easily executing arbitrary, custom EDA tool commands is the most significant advantage of the direct-Tcl approach. As a Tcl interpreter embedded in an EDA tool executes the build system code, a user can call custom commands almost anywhere.
The direct-Tcl approach is not free of drawbacks. Some of them are inherent, and others result from the fact that direct-Tcl build systems development was stopped at some stage by their maintainers.
The first inherent drawback is that Tcl is considered an unfriendly language compared to, for example, Python. The argument is often raised by junior hardware engineers and has a very subjective nature. However, Tcl was chosen as the EDA scripting language because it was very popular when EDA tools were developed (Chen et al., 2001). The second inherent drawback is that Tcl is not a popular language nowadays. Although there are no physical obstacles to change this trend, it is doubtful that suddenly more developers will start using Tcl instead of, for example, Python as an interpreted language. This is why this drawback is also treated as inherent.
The second group of direct-Tcl build systems drawbacks results from stopping development at a certain point. There is no direct-Tcl build system that can be used to examine the limits and bottlenecks of this idea. For example, most of the direct-Tcl build systems support only one EDA tool. OSVVM-Scripts is an exception. However, it can be used only for simulation. It does not support any synthesis or implementation tools. More example drawbacks common to the direct-Tcl build systems include:
@@ 2.1. Direct-Tcl approach
    Lack or limited capability of automatic cores, targets, and testbenches detection.
@@
The HBS was designed and implemented to examine the limits and bottlenecks of the direct-Tcl approach. To the author’s knowledge, there is no direct-Tcl approach hardware build system, supporting multiple EDA tools, with features and capabilities similar to those provided by the HBS.
2.2. Indirect-abstract approach
The indirect-abstract approach is much more popular than the direct-Tcl approach. In this approach, the user declares various properties and configurations rather than writing Tcl code or calling EDA tools’ Tcl commands directly. The main characteristic of the indirect-abstract approach is the execution of the build system code before the actual EDA tool is run. The result of the build system execution is a Tcl, Makefile, or a shell script, which is only later run by the specified EDA tool.
The most popular hardware build system utilizing the indirect-abstract approach is probably FuseSoc (Kindgren, 2019, 2025c). Other popular build systems include SiliconCompiler (Olofsson et al., 2022; SiliconCompiler Project, 2025), HAMMER (Liew et al., 2022; Berkeley Architecture Research, 2025), Hog (Biesuz et al., 2021; Gonnella et al., 2025), hdlmake (CERN Open Hardware Repository, 2025), orbit (Ruskin, 2025), bender (PULP Platform, 2025), bazel_rules_hdl (The Hardware Description Languages GitHub Project, 2025), flgen (Ishitani, 2025), and Blockwork (Blockwork EDA, 2025).
The indirect-abstract hardware build systems can be subdivided into the following three subcategories based on the language and syntax they utilize:
  1. (1)
    Declarative - the design build flow is described in a fully declarative way utilizing markup or data serialization languages, for example, YAML. An example of the declarative indirect-abstract build system is FuseSoc.
  2. (2)
    Programmable - the design build flow is described using a fully functional programming language, for example, Python. An example of the programmable indirect-abstract build system is SiliconCompiler.
  3. (3)
    Mixed - the design build flow is partially described using the declarative way and partially programmable. An example of the mixed indirect-abstract build system is Hog.
The language and syntax used for describing build flow vary between different build systems. Some use simple data serialization formats such as YAML (for example, FuseSoc or bender) or TOML (for example, orbit). Some come with a domain-specific language (for example, flgen), or use Python (for example, SiliconCompiler, HAMMER) and Makefiles (for example, Hdlmake). Finally, some use a mix of more than one (for example, Hog). Especially using Python as a programming language is powerful, as users can use all functionalities that come with the language.
The most significant advantage of the indirect-abstract approach is the low entry threshold. The syntax of languages such as YAML, TOML, or Python is simple, which results in high legibility. Even junior engineers without extra training can understand it. Listing 1 presents a very basic core definition in the FuseSoc build system. The first line specifies the version of the description format. The “name” key specifies the vendor, library, name, and version (VLNV) identifier specified in the IP-XACT User Guide (Accellera IP-XACT Working Group, 2018). The remaining part is relatively self-explanatory.
@@ 2.2. Indirect-abstract approach
The indirect-abstract approach is based on the concept that the interface to EDA tools is hidden under some common abstraction layer. Whether this is an advantage or a disadvantage is debatable. Providing a common abstract layer means that users do not have to learn the interface of multiple EDA tools. Moreover, in theory, changing the EDA tool for the compilation process should be limited to changing the EDA tool declaration solely. However, this is somewhat of a utopian view. As the practice shows, the common abstraction layer leads to the following common difficulties:
@@
    Missing feature implementation for some EDA tools. Let us suppose the abstraction layer declares feature F, and the build system claims support for tools T1 and T2. Although T1 might provide the implementation for F, T2 might be missing the implementation for this particular feature. In such a case, the user must add support for this feature to the build system. This task might not be trivial. The user must understand the internals of the build system, prepare a pull request, wait for a review, and potentially reiterate the implementation. Yet another approach is to modify automatically generated build scripts before a compilation. However, this is against the whole idea of build systems. A direct-Tcl approach is free of these difficulties because the user can directly insert custom Tcl commands. There are no intermediate abstraction layers.
@@
All indirect-abstract hardware build systems allow “injecting” arbitrary custom Tcl commands into the automatically generated scripts. However, these commands can not be inserted into arbitrary places. Moreover, inserting a single-line Tcl command often requires more than one line of code or even multiple files. This can be treated as proof that the indirect-abstract approach is not a golden bullet. The idea of a common abstraction layer has significant drawbacks that no one has yet found a way to fix.
Making a truly abstract hardware build system is highly unlikely in the near future. The main problem with the common abstraction layer is the fact that EDA tools were implemented before such a layer was proposed. Declaring a common interface for existing tools is a much harder task than implementing tools that conform to a formal interface.
3. Why Hardware Build System was implemented
Characterizing and comparing direct-Tcl and indirect-abstract approaches for hardware build systems theoretically did not require HBS implementation. However, none of the existing direct-Tcl build systems was functionally close to the existing indirect-abstract build systems. In particular, none of the existing direct-Tcl build systems offered the following functionalities:
@@ 3. Why Hardware Build System was implemented
    listing automatically detected cores and targets,
@@
    automatic testbench targets detection,
  4. (4)
    parallel testbench running,
  5. (5)
    dependency graph generation.
The author wanted to implement a direct-Tcl build system that would functionally adhere to the existing indirect-abstract build systems. It had to be empirically verified if it is possible to have features of indirect-abstract build systems in a direct-Tcl build system.
The implemented build system (HBS) serves as a representative of the direct-Tcl approach for the comparison with indirect-abstract build systems.
@@ 4. Hardware Build System requirements
The main goal of the HBS implementation was to push the direct-Tcl approach to its limits, examine its bottlenecks, and evaluate the concept of direct-Tcl build system by comparing it with indirect-abstract approaches. However, it does not mean that HBS is solely a prototype and was designed without thoughtful requirements. This section provides a summary of HBS major requirements with extra explanation.
@@ 4.1. Simplicity and ease of use
The primary goal of the hardware build system proposed in the paper is simplicity. Contemporary hardware build systems utilizing the indirect-abstract approach often offer simplicity only at the initial stage. However, as the project’s complexity grows, so does the complexity of using the build system.
@@ 4.2. Minimal user interface
Xia et al. (Xia et al., 2013) report that even up to 21.35% of the bugs in software build systems belong to the external interface category. This means that one-fifth of build system bugs result from incorrect usage of the user interface or different behavior of the build system with different tools or platforms.
Fixing build system external interface bugs requires applying patches in the build system itself. Such a task is relatively time-consuming and might not be simple to implement. The build system user must report the problem to the build system developers or fix the bug themselves. However, this implies that the build system user is distracted from their own work. Moreover, the fix is not applied immediately. This means that the user must also find a temporary workaround. This, in turn, leads to a technical debt (Besker et al., 2019, 2018).
A minimal user interface has one more advantage. It allows developers to customize build systems to the project’s needs. Each project is in some way unique and might require a particular approach to solve problems related to the design build. Moreover, projects usually evolve, for example, due to customer demand. Dynamic tailoring to the project requirements is a desired feature of modern build systems (Maudoux and Mens, 2018). Similar observations were made by Neitsh et al. in (Neitsch et al., 2012). Authors report that concerns from the application and implementation domains may “leak” into the build model. The build system should be adaptable enough to provide ways to respond to these concerns. Too rigid user interfaces limit such capabilities.
@@ 4.3. Low number of dependencies
Xiao et al. (Xiao et al., 2022) report that dependency management is one of the top three factors contributing to the introduction of technical debt in build systems. Such a finding is not surprising. Dependencies often break in unexpected moments. In such a case, long-term quality is sacrificed for the short-term goals. Getting a project to compile successfully is more important than fixing broken build system dependencies.
One of the goals of the HBS is to minimize the probability of user distraction caused by broken dependencies. This implies that the number of dependencies should be kept as low as possible.